OpenStudy (anonymous):

in the graph shown what is the y-coordinate of point B if the slope of line AB is 5?

OpenStudy (anonymous):

need graph

OpenStudy (anonymous):

how can i draw a graph up here lol

OpenStudy (anonymous):

The is a draw button under the reply box. Also, you can take a picture and attach it using the button below as well

OpenStudy (anonymous):



OpenStudy (anonymous):



OpenStudy (anonymous):

\[m=\frac{ \Delta y}{ \Delta x}=\frac{ y _{2}-y _{1} }{x _{2}-x _{1} }\]

OpenStudy (anonymous):

Plug in what we know and solve for y at point B

OpenStudy (anonymous):

i dont understand

OpenStudy (anonymous):

\[5=\frac{ y-4 }{ 3-2 }\]

OpenStudy (anonymous):

so is it 13

OpenStudy (anonymous):

no

OpenStudy (anonymous):

so is it 13-4=9 and the i multiplied 15-2 and got 13

OpenStudy (anonymous):

Do you see it now?

OpenStudy (anonymous):

i need more help lol

OpenStudy (anonymous):

on a different one

OpenStudy (anonymous):

So, you understand what happend with that one?

OpenStudy (anonymous):

A newborn infant usually triples its birth weight by its first birthday.Baby Carter weighed 28 pounds on his first birthday.If this were triple his birth weight (b) which equation could be used to determine his birth weight?

OpenStudy (anonymous):

Yes sir i understand the last question i had asked

OpenStudy (anonymous):

\[28=3x\]

OpenStudy (anonymous):

\[x=28/3\]

OpenStudy (anonymous):

lbs

OpenStudy (anonymous):

These are the equations shown a)b/3=28 b)b-3=28 c)3b=28 d)b+3=28 e)3(b+3)=28

OpenStudy (anonymous):

is it b/3=28

OpenStudy (anonymous):

It's c. just replace x with b

OpenStudy (anonymous):

try to explain how you did that

OpenStudy (anonymous):

We know that at 1 year a baby weighs 3 times it's birth weight. So, 3*birth weight= weight at 1year